A 10Mg truck rolling at a constant speed of 30 kph along a horizontal rail comes to an incline of slope sin 0.1. Calculate how far it will coil up the slope before coming to a halt, (a) ignoring all resistive forces, and (b) assuming the rolling vesistance is 5% of the normal reaction of the truck

1
Expert's answer
2022-02-23T08:32:59-0500

a)

"\frac{mv^2}2=mgh,"

"h=\frac{v^2}{2g},"

"l=\frac h{0.1}=35 .4~m,"

b)

"\frac{mv^2}2=mgh+Fl,"

"\frac{mv^2}2=0.1mgl+0.05mg\frac{3\sqrt{11}}{10}l,"

"l=23.7~m."